# TagRAG: Tag-guided Hierarchical Knowledge Graph Retrieval-Augmented Generation

### v1

@@ Abstract.
Retrieval-Augmented Generation enhances language models by retrieving external knowledge to support informed and grounded responses. However, traditional RAG methods rely on fragment-level retrieval, limiting their ability to address query-focused summarization queries. GraphRAG introduces a graph-based paradigm for global knowledge reasoning, yet suffers from inefficiencies in information extraction, costly resource consumption, and poor adaptability to incremental updates. To overcome these limitations, we propose TagRAG, a tag-guided hierarchical knowledge graph RAG framework designed for efficient global reasoning and scalable graph maintenance. TagRAG introduces two key components: (1) Tag Knowledge Graph Construction, which extracts object tags and their relationships from documents and organizes them into hierarchical domain tag chains for structured knowledge representation, and (2) Tag-Guided Retrieval-Augmented Generation, which retrieves domain-centric tag chains to localize and synthesize relevant knowledge during inference. This design significantly adapts to smaller language models, improves retrieval granularity, and supports efficient knowledge increment. Extensive experiments on UltraDomain datasets spanning Agriculture, Computer Science, Law, and cross-domain settings demonstrate that TagRAG achieves an average win rate of 95.41% against baselines while maintaining about 14.6x construction and 1.9x retrieval efficiency compared with GraphRAG.
@@
Retrieval-Augmented Generation, Knowledge Graph, Question Answering, Summarization, Knowledge Increment
copyright: noneconference: ; Preprint; 2025ccs: Computing methodologies Discourse, dialogue and pragmaticsccs: Information systems Question answering
1. Introduction
@@
Retrieval-Augmented Generation (RAG) (Lewis et al., 2020) is a framework that enhances the output of language models by retrieving relevant documents from an external knowledge source and conditioning the generation process on both the input query and the retrieved content, enabling more informed and factually grounded responses (Fan et al., 2024). In the era of Large Language Models (LLMs), RAG has become the most important technology to help them land effectively in the fields of medicine (Zhao et al., 2025), law (Wiratunga et al., 2024), finance (Barry et al., 2025), etc. However, constrained by its fragment-oriented retrieval, the RAG framework focuses on matching information within a limited span, leading to its lack of a global view to cope with highly abstract questions (Peng et al., 2024).
Figure 1. Inefficient graph construction and reasoning.
@@
The emergence of GraphRAG (Edge et al., 2024) solves this problem by extracting entities from documents, dividing knowledge communities, and generating community summaries, thereby refining global information. The new paradigm of introducing knowledge summarization into RAG at the graph level makes it realistic for LLMs to generate responses from a global view. However, GraphRAG still suffers from the drawbacks of inefficient information extraction and expensive resource calls. To solve these problems, some methods, such as LightRAG (Guo et al., 2024) and KET-RAG (Huang et al., 2025b), simplify the knowledge graph structure, avoiding complicated community divisions and reducing construction costs significantly. On the other hand, for improving the efficiency of inference, methods like MiniRAG (Fan et al., 2025) and FG-RAG (Hong et al., 2025) retrieve relevant subgraphs to generate query-aware fine-grained answers. However, these approaches are divorced from a global perspective. Although it is possible to inject the intrinsic knowledge of LLMs into the entity in the process of building the graph, it is difficult to compensate for the comprehensive understanding of the complete resource library.
@@
  • (RQ1) How to enhance the global knowledge graph representation and reasoning with low resource consumption? Existing Graph RAG methods heavily rely on large-scale LLMs (e.g., GPT-4o (Hurst et al., 2024)) to handle construction, retrieval, and generation, hindering migration to smaller models due to limited capacity. These high-intensity model calls also incur substantial resource costs and constrain deployment flexibility. Therefore, decoupling LLMs’ end-to-end dominance becomes essential for practical and scalable applications.
  • (RQ2) How to achieve robust and efficient knowledge increment for large-scale Graph RAG? Graph RAG employs a bottom-up construction paradigm that involves entity/relationship extraction followed by community summary generation. However, this architecture struggles with incremental updates, often requiring costly full-graph reconstruction. Consequently, developing an efficient incremental construction mechanism with minimal overhead, while ensuring seamless integration with existing knowledge structures, becomes crucial.
To tackle these challenges, we propose TagRAG, a tag-guided hierarchical knowledge graph retrieval-augmented generation framework, comprising Tag Knowledge Graph Construction and Tag-guided Retrieval-Augmented Generation. In the construction stage, TagRAG constructs a Tag Knowledge Graph by extracting object tag keywords and their relationships from documents. These object tags are linked to predefined root domain tags, forming hierarchical domain tag chains. Knowledge from object tags connected to domain tags is summarized and fused for information integration. During inference, domain-centric tags are retrieved to localize knowledge. The hierarchical tag chains are collected to integrate synthesized information, which is then fused into global answers via tag-guided retrieval-augmented generation.
In summary, the contributions are as follows.
@@
2. Related Work
Graph-based RAG strives to build graphs characterized by high-level knowledge representations and generate comprehensive responses through global search. RAPTOR (Sarthi et al., 2024) builds a hierarchical tree through clustering and summarization, transforming traditional chunk-based RAG into a global structured knowledge reasoning paradigm based on summarization. GraphRAG (Edge et al., 2024) optimizes RAPTOR’s modeling of document information as a tree to build a graph structure with richer and more diverse knowledge representations. Following GraphRAG, HiRAG (Huang et al., 2025a) and ArchRAG (Wang et al., 2025) also use a hierarchical structure based on community summaries to build and retrieve knowledge graphs. However, relying on frequent LLM calls and a large number of summarizations, the efficiency and cost of GraphRAG are unbearable for real implementation. The subsequent work has been improved in the following two aspects:
  • Structural simplification for graph construction: LightRAG (Guo et al., 2024) integrates lightweight graph structures with text indexing and dual-level retrieval, enabling efficient, adaptive knowledge access in dynamic environments. KET-RAG (Huang et al., 2025b) combines a sparse knowledge graph skeleton with a lightweight text-keyword bipartite graph, enabling efficient multi-granular retrieval without building a full-scale graph.
  • Retrieval optimization for global knowledge: MiniRAG (Fan et al., 2025) unifies text and entity indexing into a semantic-aware graph and leveraging lightweight topology-enhanced search for efficient and scalable knowledge access. FG-RAG (Hong et al., 2025) expands entity coverage through context-aware graph retrieval and enhancing response specificity with query-level fine-grained summarization. PathRAG (Chen et al., 2025) extracts key relational paths through flow-based pruning and converting them into text prompts to guide LLMs toward more coherent and context-aware generation.
@@ 2. Related Work
3. Task Definition
TagRAG is dedicated to building a hierarchically explicit knowledge graph under any domain to enable powerful and efficient RAG capabilities. Given a set of documents $D=\{d_{i}\}_{i=1}^{\left|D\right|}$, key domain information are extracted to construct an object tag knowledge graph $\mathcal{G}_{o}=(\mathcal{V}_{o},\mathcal{E}_{o})$. In order to form a clear structured knowledge management, given a root domain $\hat{v}$, a hierarchical domain tag system $\mathcal{G}_{d}=(\mathcal{V}_{d},\mathcal{E}_{d})$ needs to be established. It contains domain tags $\mathcal{V}_{d}$ at different levels and hierarchical relationships $\mathcal{E}_{d}$ between tags. For knowledge coordination and fusion, the object tags should be linked to the domain tag system and a hierarchical tag knowledge graph is finally constructed:
@@ 3. Task Definition
Figure 2. The proposed TagRAG framework.
4. TagRAG
@@
4.1. Tag Knowledge Graph Construction
@@
4.1.1. Object Tag Keyword Extraction
@@
4.1.2. Domain Tag Chain Organization
@@
Each $c_{i}$ contains multiple domain tags, where each tag consists of a domain name and its description. And the superior domain tag points to the subordinate domain tag through the relationship of “has subdomain”.
@@
1Input: The list of domain tag chain $\mathcal{C}$, the root domain tag $\hat{v}$ and the existing tag knowledge graph $\mathcal{G}$.
2 Output: $\mathcal{\hat{G}}$.
3
[…]
5 foreach $c_{i}\in\mathcal{C}$ do
6 for $j$ from 22 to $\left|c_{i}\right|-1$ do
7    $p_{i,j}\leftarrow\text{get\_node}(\text{DAG},c_{i,j-1})$;
8    $n_{i,j}\leftarrow\text{get\_node}(\text{DAG},c_{i,j})$;
9    if $p_{i,j}$ is not None then
10       if $n_{i,j}$ is not None then
11          if $n_{i,j}\notin p_{i,j}.\text{children}$ then
12             $p_{i,j}.\text{add\_child}(n_{i,j})$;
13
14
15       else
16          $n_{new}\leftarrow\text{new\_node}(c_{i,j})$;
17          $p_{i,j}.\text{add\_child}(n_{new})$;
18
19
20
21
[…]
@@ 16          $n_{new}\leftarrow\text{new\_node}(c_{i,j})$;
4.1.3. Domain-centric Knowledge Fusion
@@
$\mathcal{L}\leftarrow\{v_{i},s_{i},\text{Emb}(s_{i})\}_{i=1}^{\left|V_{d}\right|},$
where $\text{Emb}(\cdot)$ is the embedding function, $v_{i}$ denotes each domain tag, and $s_{i}$ means each summary.
4.1.4. Knowledge Incremental Insertion
@@
Compared to the time-consuming full reconstruction required by GraphRAG, TagRAG’s incremental insertion mechanism has significant efficiency advantages.
4.2. Tag-guided Retrieval-Augmented Generation
Adhering to the global vision strategy, TagRAG uses tag-guided graph retrieval-augmented generation. Because of the existence of domain tags that carry global knowledge, TagRAG is more efficient than other summarization-based and walk-based methods.
4.2.1. Domain-centric Knowledge Retrieval
@@
$(V_{t}^{{}^{\prime}},S_{t}^{{}^{\prime}})=\text{Ret}_{\text{tag}}(q,\mathcal{L}),$
@@
4.2.2. Hierarchical Tag Chain Integration
@@
4.2.3. Tag Knowledge-Fused Generation
@@
5. Experimental Setup
5.1. Datasets
@@
To demonstrate the high applicability of TagRAG, following standard evaluation on Graph-based RAG methods (Edge et al., 2024; Guo et al., 2024; Chen et al., 2025), we chose four corpus from the comprehensive UltraDomain (Qian et al., 2025) benchmark, Agriculture, CS, Legal and Mix. The size of each dataset ranges from 600,000 to 5,000,000 tokens. Following LightRAG, we used GPT-4o-mini to generate 125 global questions for each dataset, covering different domains and different tasks. The details of the datasets can be seen in Table 1. The question generation prompt can be seen in Appendix.

### v2

@@ Abstract
Retrieval-Augmented Generation enhances language models by retrieving external knowledge to support informed and grounded responses. However, traditional RAG methods rely on fragment-level retrieval, limiting their ability to address query-focused summarization queries. GraphRAG introduces a graph-based paradigm for global knowledge reasoning, yet suffers from inefficiencies in information extraction, costly resource consumption, and poor adaptability to incremental updates. To overcome these limitations, we propose TagRAG, a tag-guided hierarchical knowledge graph RAG framework designed for efficient global reasoning and scalable graph maintenance. TagRAG introduces two key components: (1) Tag Knowledge Graph Construction, which extracts object tags and their relationships from documents and organizes them into hierarchical domain tag chains for structured knowledge representation, and (2) Tag-Guided Retrieval-Augmented Generation, which retrieves domain-centric tag chains to localize and synthesize relevant knowledge during inference. This design significantly adapts to smaller language models, improves retrieval granularity, and supports efficient knowledge increment. Extensive experiments on UltraDomain datasets spanning Agriculture, Computer Science, Law, and cross-domain settings demonstrate that TagRAG achieves an average winning rate of 78.36% against baselines while maintaining about 14.6x construction and 1.9x retrieval efficiency compared with GraphRAG.
@@
TagRAG: Tag-guided Hierarchical Knowledge Graph Retrieval-Augmented Generation
Wenbiao Tao1, Xinyuan Li1, Yunshi Lan1, Weining Qian1 1East China Normal University Correspondence: yslan@dase.ecnu.edu.cn
1 Introduction
@@
Retrieval-Augmented Generation (RAG) Lewis et al. (2020) is a framework that enhances the output of language models by retrieving relevant documents from an external knowledge source and conditioning the generation process on both the input query and the retrieved content, enabling more informed and factually grounded responses Fan et al. (2024). For Large Language Models (LLMs), RAG has become the most important technology to help them land effectively in the fields of medicine Zhao et al. (2025), law Wiratunga et al. (2024), finance Barry et al. (2025), etc.
Figure 1: Inefficient graph construction and reasoning.
@@
The emergence of GraphRAG Edge et al. (2024) solves this problem by extracting entities from documents, dividing knowledge communities, and generating community summaries, thereby refining global information. The new paradigm of introducing knowledge summarization into RAG at the graph level makes it realistic for LLMs to generate responses from a global view. However, GraphRAG still suffers from the drawbacks of inefficient information extraction and expensive resource calls. To solve these problems, some methods, such as LightRAG Guo et al. (2025) and KET-RAG Huang et al. (2025b), simplify the knowledge graph structure, avoiding complicated community divisions and reducing construction costs significantly. Besides, for improving the efficiency of inference, methods like MiniRAG Fan et al. (2025) and FG-RAG Hong et al. (2025) retrieve relevant subgraphs to generate query-aware fine-grained answers. However, these approaches are divorced from a global perspective. Although it is possible to inject the intrinsic knowledge of LLMs into the entity in the process of building the graph, it is difficult to compensate for the comprehensive understanding of the complete resource library.
@@
  • (RQ1) How to enhance the global knowledge graph representation and reasoning with low resource consumption? Existing Graph RAG methods heavily rely on large-scale LLMs (e.g., GPT-4o Hurst et al. (2024)) to handle construction, retrieval, and generation, hindering migration to smaller models due to limited capacity. These high-intensity model calls also incur substantial resource costs and constrain deployment flexibility. Therefore, decoupling LLMs’ end-to-end dominance becomes essential for practical and scalable applications.
  • (RQ2) How to achieve robust and efficient knowledge increment for large-scale Graph RAG? GraphRAG employs a bottom-up construction paradigm that involves entity/relationship extraction followed by community summary generation. However, this architecture struggles with incremental updates, requiring costly full-graph reconstruction. Consequently, it is crucial to develop an efficient incremental construction mechanism that minimizes overhead and integrates seamlessly with existing knowledge structures.
To tackle these challenges, we propose TagRAG, a tag-guided hierarchical knowledge graph retrieval-augmented generation framework, comprising Tag Knowledge Graph Construction and Tag-guided Retrieval-Augmented Generation. In the construction stage, TagRAG constructs a Tag Knowledge Graph by extracting object tag keywords and their relationships from documents. These object tags are linked to predefined root domain tags, forming hierarchical domain tag chains. Knowledge from object tags connected to domain tags is summarized and fused for information integration. During inference, domain-centric tags are retrieved to localize knowledge. The hierarchical tag chains are collected to integrate synthesized information, which is then fused into global answers via tag-guided retrieval-augmented generation. In summary, the contributions are as follows:
@@
2 Related Work
Graph-based RAG Peng et al. (2024) builds graphs characterized by high-level knowledge representations and generate responses through global search. RAPTOR Sarthi et al. (2024) introduces hierarchical tree structures via clustering and summarization, shifting RAG from chunk-based retrieval to structured reasoning. GraphRAG Edge et al. (2024) further enriches this modeling with diverse graph representations, followed by HiRAG Huang et al. (2025a) and ArchRAG Wang et al. (2025), which utilize hierarchical community summaries. However, the heavy reliance on frequent LLM calls and extensive summarization makes GraphRAG prohibitively expensive and inefficient for practical deployment.
The subsequent work has been improved in the following two aspects: (1)Structural simplification for graph construction. LightRAG Guo et al. (2025) integrates lightweight graphs with text indexing and dual-level retrieval for efficient, adaptive knowledge access. KET-RAG Huang et al. (2025b) combines a sparse knowledge graph skeleton with a text-keyword bipartite graph, enabling multi-granular retrieval without building a full-scale graph. (2)Retrieval optimization for global knowledge. MiniRAG Fan et al. (2025) unifies text and entity indexing into a semantic-aware graph, leveraging lightweight topology-enhanced search for efficient knowledge access. FG-RAG Hong et al. (2025) extends entity coverage via context-aware graph retrieval and improves response specificity with query-level fine-grained summarization. PathRAG Chen et al. (2025) extracts key relational paths and converts them into text prompts, guiding LLMs toward more coherent and context-aware generation.
@@ 2 Related Work
3 Task Definition
TagRAG is dedicated to building a hierarchically explicit knowledge graph under any domain to enable powerful and efficient RAG capabilities. Given a set of documents $D=\{d_{i}\}_{i=1}^{\left|D\right|}$, key domain information are extracted to construct an object tag knowledge graph $\mathcal{G}_{o}=(\mathcal{V}_{o},\mathcal{E}_{o})$. In order to form a clear structured knowledge management, given a root domain $\hat{v}$, a hierarchical domain tag system $\mathcal{G}_{d}=(\mathcal{V}_{d},\mathcal{E}_{d})$ needs to be established. It contains domain tags $\mathcal{V}_{d}$ at different levels and hierarchical relationships $\mathcal{E}_{d}$ between tags. For knowledge coordination and fusion, the object tags should be linked to the domain tag system and a hierarchical tag knowledge graph is constructed:
@@ 3 Task Definition
Figure 2: The proposed TagRAG framework.
4 TagRAG
@@
4.1 Tag Knowledge Graph Construction
@@
4.1.1 Object Tag Keyword Extraction
@@
4.1.2 Domain Tag Chain Organization
@@
Each $c_{i}$ contains multiple domain tags, each consisting of a domain name and its description. Superior domain tag points to the subordinate domain tag through the relationship of "has subdomain".
@@
Input: The list of domain tag chain $\mathcal{C}$, the root domain tag $\hat{v}$ and the existing tag knowledge graph $\mathcal{G}$.
Output: $\mathcal{\hat{G}}$.
$\text{DAG}\leftarrow\{\hat{v}\}$;
foreach $c_{i}\in\mathcal{C}$ do
 for $j$ from 22 to $\left|c_{i}\right|-1$ do
    $p_{i,j}\leftarrow\text{get\_node}(\text{DAG},c_{i,j-1})$;
    $n_{i,j}\leftarrow\text{get\_node}(\text{DAG},c_{i,j})$;
    if $p_{i,j}$ is not None then
       if $n_{i,j}$ is not None then
          if $n_{i,j}\notin p_{i,j}.\text{children}$ then
             $p_{i,j}.\text{add\_child}(n_{i,j})$;
       else
          $n_{new}\leftarrow\text{new\_node}(c_{i,j})$;
          $p_{i,j}.\text{add\_child}(n_{new})$;
$\mathcal{\hat{G}}\leftarrow\text{merge\_graph}(\mathcal{G},\text{DAG})$;
@@
4.1.3 Domain-centric Knowledge Fusion
@@
$\mathcal{K}\leftarrow\{v_{i},s_{i},\text{Emb}(s_{i})\}_{i=1}^{\left|V_{d}\right|},$
where $\text{Emb}(\cdot)$ is the embedding function, $v_{i}$ and $s_{i}$ denote a domain tag and its summary, respectively.
4.1.4 Knowledge Incremental Insertion
@@
Compared to the time-consuming full reconstruction of GraphRAG, TagRAG’s incremental insertion mechanism is significantly more efficient.
4.2 Tag-guided Retrieval-Augmented Generation
Following the global vision strategy, TagRAG uses tag-guided graph retrieval-augmented generation. As the existence of domain tags that carry global knowledge, TagRAG is more efficient than other summarization-based and walk-based methods.
4.2.1 Domain-centric Knowledge Retrieval
@@
$(V_{t}^{{}^{\prime}},S_{t}^{{}^{\prime}})=\text{Ret}_{\text{tag}}(q,\mathcal{K}),$
@@
4.2.2 Hierarchical Tag Chain Integration
@@
4.2.3 Tag Knowledge-Fused Generation
@@
4.3 Analysis of Retrieval Complexity
GraphRAG’s inefficiency stems from its multi-stage pipeline, requiring extensive LLM calls for entity extraction and community detection (e.g., Leiden algorithm), leading to high computational costs and frequent full graph reconstructions for dynamic data. In contrast, TagRAG’s domain tag chains enable hierarchical information aggregation during graph building, replacing iterative community partitioning with linear chain processing. During inference, TagRAG utilizes vector matching and tag chain linking, avoiding GraphRAG’s costly full graph traversal. Both stages demonstrate TagRAG’s superior efficiency over GraphRAG.
5 Experimental Setup
5.1 Datasets and Baselines
@@
To demonstrate the high applicability of TagRAG, following standard evaluation on Graph-based RAG methods Edge et al. (2024); Guo et al. (2025); Chen et al. (2025), we chose four corpus from the comprehensive UltraDomain Qian et al. (2025) benchmark, Agriculture, CS, Legal and Mix. Following LightRAG, we used GPT-4o-mini to generate 125 global questions for each dataset, covering different domains and different tasks. The dataset details and question generation prompt can be seen in Appendix A.2 and C.
To validate the performance and efficiency of global question answering, we compare TagRAG with two types of baselines. (1) Zero-shot LLM Generation: We call Qwen3-4B Team (2025), Qwen3-30B-A3B Team (2025) and Llama-3.3-70B-Instruct Grattafiori et al. (2024) to directly answer the questions. (2) Retrieval-Augmented Generation: NaiveRAG Lewis et al. (2020) focuses on local context through dynamic document retrieval. GraphRAG Edge et al. (2024) utilizes entity graphs and community summaries for global knowledge synthesis. LightRAG Guo et al. (2025) is a lightweight baseline that integrates graph structures with dual-level text indexing for fast and real-time updates. MiniRAG Fan et al. (2025) achieves high efficiency on small language models via semantic-aware graph indexing and lightweight topology-based retrieval. Detailed descriptions are shown in Appendix A.3.
5.2 Evaluation Metrics
Four evaluation metrics were used to assess the performance of the comparison methods: Comprehensiveness: How much detail does the answer provide to cover all aspects and details of the question? Diversity: How varied and rich is the answer in providing different perspectives and insights on the question? Empowerment: How well does the answer help the reader understand and make informed judgments about the topic? Overall: Which answer is better overall?
We utilize the powerful model GPT-4o-mini, gemini-2.5-pro and claude-sonnet-4.5-20250929 to determine the winner for each of the two comparison methods based on the above metrics. For each dataset, we also exchange the order of the results of the two comparison methods to avoid position bias. Finally, we report the average of 3*2=6 evaluation results. The evaluation prompt can be seen in Appendix C.
5.3 Implementation Details
We use Qwen3-4B Team (2025) as the backbone to conduct the experiments without thinking. bge-large-en-v1.5 Xiao et al. (2023) is employed to embed questions and documents. The chunk size and overlap size are 1200 and 100, respectively. The vector database used in this work is nano-vectordb  [^1] https://github.com/gusye1234/nano-vectordb. The top-k number for Domain-centric Knowledge Retrieval is 3. More detailed experimental settings can be found in Appendix A.
Agri CS Legal Mix Avg
[…]
v.s. Qwen3-4B
Comprehensiveness 85.9 69.6 70.3 58.8 71.2
Diversity 78.9 64.4 62.0 53.9 64.8
Empowerment 80.7 57.5 48.9 45.7 58.2
Overall 83.2 59.3 53.7 49.3 61.4
v.s. Qwen3-30B-A3B
Comprehensiveness 84.8 68.8 72.7 59.6 71.5
Diversity 79.5 65.1 62.1 59.6 66.6
Empowerment 76.4 53.2 45.6 40.4 53.9
Overall 80.3 55.5 51.6 44.7 58.0
v.s. Llama-3.3-70B-Instruct
Comprehensiveness 68.1 51.5 49.7 27.7 49.2
Diversity 67.5 58.5 43.1 35.9 51.2
Empowerment 56.9 43.1 27.2 22.8 37.5
Overall 65.9 49.3 39.7 27.3 45.5
[…]
v.s. NaiveRAG
Comprehensiveness 93.9 94.0 79.5 90.0 89.3
Diversity 90.4 95.5 95.2 95.2 94.1
Empowerment 91.2 88.3 74.9 88.1 85.6
Overall 93.5 89.7 72.3 87.9 85.8
v.s. GraphRAG
Comprehensiveness 89.3 85.3 60.7 67.1 75.6
Diversity 88.5 85.5 72.0 74.9 80.2
Empowerment 90.8 81.6 63.2 67.9 75.9
Overall 90.9 81.9 62.3 67.7 75.7
v.s. LightRAG
Comprehensiveness 94.4 85.1 80.3 88.1 87.0
Diversity 94.9 88.1 88.5 92.3 91.0
Empowerment 95.5 85.7 82.0 91.6 88.7
Overall 94.4 84.0 79.9 89.7 87.0
v.s. MiniRAG
Comprehensiveness 76.0 55.2 56.8 64.1 63.0
Diversity 80.7 64.7 70.4 77.1 73.2
Empowerment 79.1 59.7 59.5 67.3 66.4
Overall 78.3 58.4 57.5 65.3 64.9
Table 1: Main results: winning rates (%) of TagRAG v.s. baselines with Qwen3-4B across four datasets.
Refer to caption Refer to caption
(a) On UltraDomain Agriculture (b) On UltraDomain CS
Refer to caption Refer to caption
(c) On UltraDomain Legal (d) On UltraDomain Mix
Figure 3: Performance-efficiency analysis: comparative winning rates, graph construction time and inference time results of TagRAG and baselines across four datasets. The larger the bubble and the closer to the lower left corner, the better the method.
Refer to caption Refer to caption
(a) Comprehensiveness (b) Diversity
Refer to caption Refer to caption
(c) Empowerment (d) Overall
Figure 4: Incremental analysis: winning rates (%) of TagRAG v.s. baselines with Qwen3-4B across four datasets in 10 rounds. The red horizontal line represents equilibrium and values above it indicate TagRAG’s advantage.
Agri CS Legal Mix
v.s. w/o chain
Comprehensiveness 87.1 90.5 85.2 78.9
Diversity 84.7 87.3 85.6 76.3
Empowerment 86.8 84.5 77.7 72.8
Overall 87.2 87.5 80.1 75.2
v.s. w/o fusion
Comprehensiveness 97.3 95.5 95.5 85.9
Diversity 96.4 94.1 94.5 89.1
Empowerment 96.9 87.1 85.1 76.9
Overall 96.9 89.7 88.0 78.4
Table 2: Ablation study: winning rates (%) of TagRAG v.s. w/o chain and w/o fusion with Qwen3-4B across four datasets.
6 Results and Analysis
6.1 Main Results
Table 1 evaluates the winning rate of TagRAG against NaiveRAG, GraphRAG, LightRAG, and MiniRAG on four datasets: UltraDomain Agriculture, CS, Legal, and Mix. We can draw the following conclusions: (1)TagRAG leads in performance, achieving an average winning percentage of 82.85% compared with NaiveRAG, GraphRAG, and LightRAG. Even against MiniRAG, the strongest baseline, TagRAG maintains a dominant average winning rate, with MiniRAG retaining only 35.125%. (2)TagRAG excels at Diversity, leveraging its domain knowledge integration to access broader information for generation. Even when compared to GraphRAG, which traverses all communities, TagRAG maintains a clear advantage, suggesting that fusing knowledge via domain chains is superior to community-summary or path-discovery approaches. (3)TagRAG expands the model capability, because it defeats the LLM with 7 times the parameters by virtue of the fused domain chain information. Even facing the 70B model, TagRAG can compete with it. This demonstrates that TagRAG does not rely solely on the inherent capabilities of the language model, and that the tag chain mechanism effectively broadens the boundaries of global reasoning.
6.2 Performance-efficiency Analysis
In Figure 3, we show the performance, graph construction time and inference time of different methods. TagRAG, which presents the largest bubbles, embodies the strongest performance, and possesses low graph construction time and inference time. GraphRAG not only exhibits weak performance, but also consumes extremely high construction and inference costs. MiniRAG has the next strongest performance, but its graph construction time is far behind TagRAG, due to its ultimate entity and relationship extraction. Despite the excellent time overhead with lightweight extraction, the performance of LightRAG is the worst, subject to the gap of high and low level semantics. Taken together, thanks to the advanced domain chain indexing and knowledge integration, TagRAG has achieved the strongest performance-efficiency results.
6.3 Ablation Study
We conduct ablation experiments on the four datasets by comparing TagRAG with TagRAG w/o chain and w/o fusion. TagRAG w/o chain simply adds the retrieved fused domain-centric knowledge to the context without involving the information in the associated domain chain. And TagRAG w/o fusion not only discards the knowledge of the domain chain, but also cuts off the object tag information connected to the retrieved domain tag. It relies entirely on the description of the domain tags to answer the question. It can be clearly found that (1) the absence of these two components makes the answers generated by the model far inferior to the full TagRAG and (2) retaining fused domain-centric knowledge is better than simply using the description of domain tags for generation.
6.4 Incremental Analysis
To validate that our proposed method can accommodate incremental scenarios, we incrementally construct a knowledge graph from 10 documents in the UltraDomain CS dataset. After each round of increment, we generate diverse questions involving all current documents to test the global reasoning capabilities of different methods. In Figure 4, Tag RAG has a steady lead over Naive RAG, Graph RAG and Light RAG, with average winning rates reaching over 80%. Even though there is a close tie with TagRAG in the first round, MiniRAG loses as the documents increase. This proves that TagRAG has a stable global inference capability in incremental scenarios with multiple rounds, as the mechanism of tag chain fusion naturally has the advantage of infinite scaling.
v.s. NaiveRAG v.s. GraphRAG v.s. LightRAG v.s. MiniRAG
large base small large base small large base small large base small
Comprehensiveness 94.0 93.7 93.1 85.3 86.0 86.1 85.1 81.5 82.3 55.2 57.5 59.9
Diversity 95.5 95.6 96.3 85.5 85.6 87.2 88.1 87.3 86.8 64.7 69.2 66.0
Empowerment 88.3 90.9 90.5 81.6 84.4 85.2 85.7 84.9 85.7 59.7 62.8 64.5
Overall 89.7 89.9 91.1 81.9 82.8 85.1 84.0 82.5 83.3 58.4 60.7 62.8
Table 3: Retriever adaption analysis: winning rates (%) of TagRAG v.s. baselines with Qwen3-4B on CS, equiped with different retrievers of bge-large-en-v1.5, bge-base-en-v1.5 and bge-small-en-v1.5.
6.5 Retriever adaption analysis
To demonstrate that our performance does not solely depend on the effect of a powerful searcher, we perform a retriever adaption analysis. As shown in Table 3, retrievers of different sizes are applied to the comparison. Even with a lower performance retriever, TagRAG still has a stable advantage in all metrics. Even in comparison with GraphRAG, the poorer retriever excites the potential of TagRAG. This suggests that domain clustering in the form of tag chains based on DAG weakens the need for precise entity retrieval and enhances the representation of global knowledge.
7 Conclusion
We propose TagRAG, a tag-guided hierarchical knowledge graph RAG framework that addresses key limitations of existing graph-based RAG systems, such as inefficient global reasoning and poor increment adaptability. By constructing domain-aware tag chains and enabling tag-guided graph retrieval, TagRAG achieves structured, fine-grained, and scalable knowledge integration. Our design reduces reliance on large language models and supports efficient construction and inference. The experimental results on the UltraDomain benchmark show that TagRAG achieves a 78.36% average winning rate over baselines while offering significant gains in construction (14.6×) and retrieval (1.9×) efficiency compared to GraphRAG.
Limitations
The indexing pipeline of TagRAG depends on LLM calls for object tag extraction and domain chain construction, which raises questions about cost, reproducibility, and robustness in fully automated scenarios. In addition, TagRAG cannot handle multimedia data such as pictures and videos, which limits its application scope.
Appendix A Detailed Experimental Settings
A.1 Experimental Environment
Configuration
CPU Intel(R) Xeon(R) Gold 6330 CPU
@ 2.00GHz
GPU NVIDIA RTX A6000 48GB
RAM 256GB
Operating System CentOS Linux 7
CUDA 12.4
Python 3.10.16
Table 4: Experimental environment.
A.2 Dataset Details
We conduct experiments on four datasets, namely Agriculture, CS, Legal and Mix. Agriculture, CS and Legal describe expertise in their respective fields, while Mix mixes knowledge from 18 fields. As shown in Table 5, the Agriculture and CS datasets are comparable in size, with Legal being the largest and Mix being the smallest.